# Leveraging configuration interaction singles for qualitative descriptions of ground and excited states: state-averaging, linear-response, and spin-projection

### v1

@@ Abstract
While configuration interaction singles (CIS) provides a computationally efficient description of excited states, it systematically overestimates excitation energies and performs poorly for strongly correlated systems, partially due to the lack of orbital relaxation and the strong ground-state bias of Hartree–Fock orbitals. To address these limitations, we present a unified variational framework that extends CIS by incorporating orbital optimization (state-specific and state-averaged), linear-response orbital relaxation via double-CIS schemes, and spin-symmetry breaking and restoration. In spin-projected state-averaged formulations, standard multistate parametrizations are no longer valid because the projection operator breaks the unitary invariance of orbital rotations and induces nonorthogonal couplings among states. By formulating a rigorous state-averaged objective in the projected subspace, we derive analytic gradients and Hessians and enable robust optimization using a trust-region augmented Hessian algorithm. Benchmark calculations show that spin projection alone significantly exacerbates the CIS overestimation in weakly correlated systems, whereas combining spin projection with state averaging or double-CI corrections substantially reduces errors, particularly for Rydberg excitations. We further demonstrate that state averaging and spin projection provide complementary and essential benefits in strongly correlated regimes, as illustrated by the bond dissociation of hydrogen fluoride.
@@ I Introduction
Accurate yet computationally affordable descriptions of electronically excited states remain a long-standing challenge in quantum chemistry. High-level wave-function methods such as equation-of-motion coupled-cluster theory with singles and doubles or triples (EOM-CCSD/CCSDT)Stanton and Bartlett (1993); Krylov (2008); Bartlett and Musiał (2007) and other related methodsChristiansen, Koch, and JÃÿrgensen (1995); Koch et al. (1997) provide reliable excitation energies for a wide range of molecular systems. However, their steep computational cost severely restricts their applicability to large molecules, extended excited-state manifolds, and the exploration of excited-state potential energy surfaces. This limitation has motivated sustained efforts to develop low-cost excited-state methods that capture essential physical effects while retaining the favorable scaling and conceptual simplicity of mean-field–based approaches.
@@
Among such methods, configuration interaction singles (CIS) represents one of the most fundamental low-cost approaches to excited states. Formally equivalent to the Tamm–Dancoff approximation (TDA) to time-dependent Hartree–Fock (TDHF), CIS provides a variationally stable and computationally efficient description of excited states in terms of single excitations from a Hartree–Fock (HF) reference. Despite these appealing features, CIS is well known to systematically overestimate excitation energies.Hirata and Head-Gordon (1999); Dreuw and Head-Gordon (2005); Subotnik (2011) This deficiency originates from two fundamental limitations: the absence of dynamical correlation and, more importantly for many classes of excitations, the lack of orbital relaxation. Because excitation energies are differences between ground- and excited-state energies, their errors arise from imbalances in correlation energy between the two states. In particular, the missing differential correlation energy is typically on the order of the correlation energy associated with the electron pairs involved in the excitation. Dynamical correlation effects can often be treated perturbatively by including double excitations with respect to CIS states, leading to the widely used CIS(D) approach.Head-Gordon et al. (1994) While treatments of dynamical correlation have a long history and are now relatively well established, the incorporation of orbital relaxation effects for excited states has attracted renewed interest in recent years.Gilbert, Besley, and Gill (2008); Kowalczyk et al. (2013); Levi, Ivanov, and Jónsson (2020); Hait and Head-Gordon (2021); Kossoski and Loos (2023); Damour et al. (2024); Tuckman and Neuscamman (2023); Kussmann et al. (2024); Selenius et al. (2024)
@@
Because HF orbitals are optimized solely for the ground state, CIS excited states are inherently biased toward the ground-state electronic structure, leading to a systematic imbalance between ground and excited states. This bias is particularly severe for excitations that involve substantial redistribution of electron density such as charge-transferSubotnik (2011); Liu and Subotnik (2014) and core excited states,Norman and Dreuw (2018); Liu et al. (2019); Oosterbaan, White, and Head-Gordon (2019); Garner and Neuscamman (2020) and for near-degenerate electronic states.Tsuchimochi and Van Voorhis (2015); Tsuchimochi (2015) Consequently, CIS, or more broadly, single-reference linear-response methods including TDDFT, also fails qualitatively in strongly correlated regimes, such as bond dissociation and conical intersections,Levine et al. (2006); Li et al. (2014); Taylor, Tozer, and Curchod (2023) where the single-reference picture breaks down.
@@
Orbital relaxation therefore plays a central role in improving CIS-based excited-state methods. One rigorous way to incorporate orbital relaxation is full state-specific orbital optimization for each excited state. This philosophy underlies excited-state mean-field theory (ESMF), introduced by Neuscamman and co-workers,Shea and Neuscamman (2018); Shea, Gwin, and Neuscamman (2020); Tran and Neuscamman (2020); Hardikar and Neuscamman (2020) which can be viewed as a state-specific, orbital-optimized CIS (SSCIS) method. In ESMF, or equivalently SSCIS, orbitals and CI coefficients are optimized simultaneously for a targeted excited state, yielding substantial improvements in excitation energies through balanced orbital relaxation especially when combined with perturbative treatments of dynamical correlation.Shea and Neuscamman (2018); Clune, Shea, and Neuscamman (2020); Clune and Neuscamman (2025) However, this approach introduces a highly nonlinear optimization problem that is variationally unstable and potentially difficult to converge in practice. Moreover, the resulting excited states are state-specific and nonorthogonal, complicating the description of multiple excited states and the evaluation of transition properties.
@@
To mitigate the strong ground-state bias inherent in HF-based excited-state methods, state averaging provides yet another route. By optimizing orbitals with respect to an average energy over several electronic states, state-averaged approaches reduce the preferential stabilization of the ground state and yield a more balanced description of excited states. State averaging has a long history in multiconfigurational self-consistent field (MCSCF) theoriesWerner and Meyer (1981); Szalay et al. (2012), and also has been combined with other methods,Kazaryan, Heuver, and Filatov (2008); Nite and Jiménez-Hoyos (2019) particularly for photochemical applications and near-degenerate electronic structures. When applied to CIS-based methods, state-averaged CIS (SACIS) should provide a natural framework for introducing orbital relaxation “on average” across multiple states, as a variationally stable alternative to state-specific orbital optimization. However, whether and to what extent this strategy mitigates the systematic overestimation of CIS excitation energies remains to be systematically assessed.
@@
In this paper, we present a unified framework for state-specific and state-averaged orbital-optimized CIS and ECIS, including their linear-response extensions based on DCIS. We derive analytic gradients and Hessians of SACIS and SAECIS, formulate efficient and robust optimization strategies using TRAH, and systematically assess the performance of these methods for both weakly and strongly correlated systems. Through benchmark calculations and detailed analyses of convergence behavior and potential energy surfaces, we aim to clarify the respective roles of orbital relaxation, state averaging, and spin projection in low-cost excited-state methods, and to delineate their regimes of applicability.
@@ II.1.1 Energy, gradients, and Hessian
If the spin-projection is neglected, this reduces exactly to the excited state mean-field theoryShea and Neuscamman (2018); Shea, Gwin, and Neuscamman (2020); Tran and Neuscamman (2020); Hardikar and Neuscamman (2020).
@@ II.1.2 Treatment of other excited states: double CIS
One of the main disadvantages of SSCIS/SSECIS is that orbitals are optimized solely with respect to a particular CIS/ECIS state, which significantly deteriorates the description of other CIS/ECIS states. For instance, when the ground state is optimized using generalized CIS/ECIS, the other solutions, which are orthogonal to each other and thus correspond to excited states, exhibit excessively high energies, leading to a substantial overestimation of excitation energies. In the ground state SSECIS, the overestimation is even more pronounced. To avoid this issue, one could perform state-specific calculations independently for different states of interest in the same spirit as $\Delta$ SCF; however, the use of different orbitals makes the optimized states non-orthogonal to each other, complicating the computation of inter-state quantities such as transition dipole moments.
@@ II.2.1 Energy, gradients, and Hessian
It should first be noted that, for multistate calculations of standard MCSCF (without $\hat{P}$) such as SA-CASSCF, one frequently employs the following parameterization:Helgaker, Jørgensen, and Olsen (2000)
@@
The second derivatives are important not only for orbital optimization but also for geometry optimization (nuclear gradients), as will be discussed in a separate paper.Tsuchimochi Again, in real orbitals,
@@ III.1 Newton method and trust-region augmented Hessian method
However, this update is often not optimal when the states are (on average) far from the minimum, where the energy functional is not well approximated by a quadratic form.Helgaker, Jørgensen, and Olsen (2000) Instead, the trust-region augmented Hessian (TRAH) algorithm diagonalizes the augmented Hessian,
@@ III.3 Point-group symmetry
In spin-projection approaches, point-group symmetry is not explicitly enforced, and the converged wave function is therefore not guaranteed to retain the correct point-group symmetry, since spin-unrestricted methods allow the breaking of spatial symmetry in addition to spin symmetry. To explicitly account for point-group symmetry within the framework of symmetry breaking and restoration, a point-group symmetry projection may be desirable.Jiménez-Hoyos et al. (2012); Jiménez-Hoyos, Rodríguez-Guzmán, and Scuseria (2013); Uejima and Ten-no (2017) However, from our experiences in many cases this is not needed; spin-projection can resurrect the lost spatial symmetry if not all. This means each of the contaminated spin states possesses a different point-group symmetry than the one the designated spin state has. Therefore, we will not perform the point-group symmetry projection, but simply assign each excited state to the dominant point-group irreducible representation (irrep).
@@ IV.1 Computational details
All CIS, ECIS, SACIS, SAECIS, DCIS, and EDCIS calculations were performed using the Gellan suite of programs.Gel For complete-active-space SCF (CASSCF) and equation-of-motion coupled-cluster with singles and doubles (EOM-CCSD), we used Q-Chem.Epifanovsky et al. (2021) The 6-31G basis set was used for hydrogen fluoride, while aug-cc-pVDZ was employed for all other molecules.
@@
For the benchmark calculations in Section IV.3, we state-averaged over a fixed manifold of $n$ states that includes prescribed numbers of roots from each irrep in the benchmark set (including the ground state). In all methods, however, point-group symmetry constraints were not enforced during the orbital optimization. Therefore, for the SA optimization of multiple states with designated symmetry, we employed DIIS and, in each macro iteration, first converged more than $n$ CIS/ECIS roots in the micro step. From these converged roots, we then identified and selected the lowest-energy solutions with the desired irreps following the protocol described in Section III.3, and used only this selected set to construct the SA orbital gradient for the subsequent DIIS update.
@@ IV.3 Benchmark calculations for excitation energies of weakly correlated systems
Finally, we emphasize that the conclusions drawn here are specific to weakly correlated regimes. We expect that the proposed methods, particularly SAECIS and EDCIS, will perform more favorably for strongly correlated systems, where orbital relaxation and near-degeneracy effects play a more prominent role. A systematic assessment in such regimes, however, requires appropriate benchmark datasets, which are currently unavailable and will therefore be pursued in future work. In the absence of such a dataset, next we use the bond dissociation of hydrogen fluoride as a stringent qualitative test case for the proposed methods in a strongly correlated regime.
@@ IV.4 Potential energy curves of hydrogen fluoride
where $|\rm(Core)\rangle=|\phi_{{\rm F}_{1s}}^{2}\phi_{{\rm F}_{2s}}^{2}\phi_{{\rm F}_{p_{x}}}^{2}\phi_{{\rm F}_{p_{y}}}^{2}\rangle$ denotes the closed-shell, non-bonding orbitals. Inspection of the ground-state wave function at large bond distances reveals that SACIS instead favors a localized-orbital description reminiscent of valence bond theory. In this limit, the reference determinant $|\Phi_{0}\rangle$ corresponds to the ionic configuration H+F-. Single excitations from $\phi_{{\rm F}_{p_{z}}}$ to $\phi_{{\rm H}_{1s}}$ within CIS then yield the correct dissociation limit in the form of a charge-transfer singlet,
@@
As shown in Figure 4(b), the absolute ground-state energy errors of SACIS relative to FCI are comparable to those of SUHF. However, the SACIS curve is more nearly parallel to the FCI reference. This is quantified by the non-parallelity error (NPE), defined as the difference between the maximum and minimum deviations from FCI, which is reduced from 13.9 mHartree for SUHF to 9.6 mHartree for SACIS, as summarized in Table 3.

### v2

@@ Abstract
While configuration interaction singles (CIS) provides a computationally efficient description of excited states, it systematically overestimates excitation energies and performs poorly for strongly correlated systems, partly due to the lack of orbital relaxation and the strong ground-state bias of Hartree–Fock orbitals. To address these limitations, we present a unified variational framework that extends CIS by incorporating orbital optimization in state-specific and state-averaged forms (SSCIS and SACIS), linear-response orbital relaxation via a double-CIS scheme (DCIS), and spin-symmetry breaking and restoration (ECIS). In spin-projected state-averaged formulations, standard multistate parametrizations are no longer valid because the projection operator breaks the unitary invariance of orbital rotations and induces nonorthogonal couplings among states. By formulating a rigorous state-averaged objective in the projected subspace, we derive analytic electronic gradients and Hessians and enable robust optimization using a trust-region augmented Hessian algorithm. Benchmark calculations show that spin projection alone significantly exacerbates the CIS overestimation in weakly correlated systems, whereas combining spin projection with state averaging or double-CIS corrections substantially reduces errors, particularly for Rydberg excitations. We further demonstrate that state averaging and spin projection provide complementary and essential benefits in strongly correlated regimes, as illustrated by the bond dissociation of hydrogen fluoride and nitrogen.
@@ I Introduction
Accurate yet computationally affordable descriptions of electronically excited states remain a long-standing challenge in quantum chemistry. High-level wave-function methods such as equation-of-motion coupled-cluster theory with singles and doubles or triples (EOM-CCSD/CCSDT)Stanton and Bartlett (1993); Krylov (2008); Bartlett and Musiał (2007) and other related methodsChristiansen et al. (1995); Koch et al. (1997) provide reliable excitation energies for a wide range of molecular systems. However, their steep computational cost severely restricts their applicability to large molecules, extended excited-state manifolds, and the exploration of excited-state potential energy surfaces. This limitation has motivated sustained efforts to develop low-cost excited-state methods that capture essential physical effects while retaining the favorable scaling and conceptual simplicity of mean-field–based approaches.
@@
Among such methods, configuration interaction singles (CIS) represents one of the most fundamental low-cost approaches to excited states. Formally equivalent to the Tamm–Dancoff approximation (TDA) to time-dependent Hartree–Fock (TDHF), CIS provides a variationally stable and computationally efficient description of excited states in terms of single excitations from a Hartree–Fock (HF) reference. Despite these appealing features, CIS is well known to systematically overestimate excitation energies.Hirata and Head-Gordon (1999); Dreuw and Head-Gordon (2005); Subotnik (2011) This deficiency originates from two fundamental limitations: the absence of dynamical correlation and, more importantly for many classes of excitations, the lack of orbital relaxation. Because excitation energies are differences between ground- and excited-state energies, their errors arise from imbalances in correlation energy between the two states. In particular, the missing differential correlation energy is typically on the order of the correlation energy associated with the electron pairs involved in the excitation. Dynamical correlation effects can often be treated perturbatively by including double excitations with respect to CIS states, leading to the widely used CIS(D) approach.Head-Gordon et al. (1994) While treatments of dynamical correlation have a long history and are now relatively well established, the incorporation of orbital relaxation effects for excited states has attracted renewed interest in recent years.Gilbert et al. (2008); Kowalczyk et al. (2013); Levi et al. (2020); Hait and Head-Gordon (2021); Kossoski and Loos (2023); Damour et al. (2024); Tuckman and Neuscamman (2023); Kussmann et al. (2024); Selenius et al. (2024)
@@
Because HF orbitals are optimized solely for the ground state, CIS excited states are inherently biased toward the ground-state electronic structure, leading to a systematic imbalance between ground and excited states. This bias is particularly severe for excitations that involve substantial redistribution of electron density, such as charge-transferSubotnik (2011); Liu and Subotnik (2014) and core-excited states,Norman and Dreuw (2018); Liu et al. (2019); Oosterbaan et al. (2019); Garner and Neuscamman (2020) as well as for near-degenerate electronic states.Tsuchimochi and Van Voorhis (2015); Tsuchimochi (2015) In addition, CIS, or more broadly, single-reference linear-response methods including TDDFT, also fail qualitatively in strongly correlated regimes, such as bond dissociation and conical intersections,Levine et al. (2006); Li et al. (2014); Taylor et al. (2023) where the single-reference picture breaks down.
@@
Orbital relaxation therefore plays a central role in improving CIS-based excited-state methods. One rigorous way to incorporate orbital relaxation is full state-specific orbital optimization for each excited state. This philosophy underlies excited-state mean-field theory (ESMF), introduced by Neuscamman and co-workers,Shea and Neuscamman (2018); Shea et al. (2020); Tran and Neuscamman (2020); Hardikar and Neuscamman (2020) which can be viewed as a state-specific, orbital-optimized CIS (SSCIS) method. In ESMF, or equivalently SSCIS, orbitals and CI coefficients are optimized simultaneously for a targeted excited state, yielding substantial improvements in excitation energies through balanced orbital relaxation especially when combined with perturbative treatments of dynamical correlation.Shea and Neuscamman (2018); Clune et al. (2020); Clune and Neuscamman (2025) However, this approach introduces a highly nonlinear optimization problem that is variationally unstable and potentially difficult to converge in practice. Moreover, the resulting excited states are state-specific and nonorthogonal, complicating the description of multiple excited states and the evaluation of transition properties.
@@
To mitigate the strong ground-state bias inherent in HF-based excited-state methods, state averaging provides yet another route. By optimizing orbitals with respect to an average energy over several electronic states, state-averaged approaches reduce the preferential stabilization of the ground state and yield a more balanced description of excited states. State averaging has a long history in multiconfigurational self-consistent field (MCSCF) theoriesWerner and Meyer (1981); Szalay et al. (2012), and also has been combined with other methods,Kazaryan et al. (2008); Nite and Jiménez-Hoyos (2019) particularly for photochemical applications and near-degenerate electronic structures. When applied to CIS-based methods, state-averaged CIS (SACIS) should provide a natural framework for introducing orbital relaxation “on average” across multiple states, as a variationally stable alternative to state-specific orbital optimization. However, whether and to what extent this strategy mitigates the systematic overestimation of CIS excitation energies remains to be systematically assessed.
@@
In this paper, we present a unified framework for state-specific and state-averaged orbital-optimized CIS and ECIS, including their linear-response extensions based on DCIS. We derive analytic gradients and Hessians of SACIS and SAECIS with respect to orbital rotations and CI coefficients, formulate efficient and robust optimization strategies using TRAH, and systematically assess the performance of these methods for both weakly and strongly correlated systems. Through benchmark calculations and detailed analyses of convergence behavior and potential energy surfaces, we aim to clarify the respective roles of orbital relaxation, state averaging, and spin projection in low-cost excited-state methods, and to delineate their regimes of applicability.
@@ II.1.1 Energy, gradients, and Hessian
If the spin-projection is neglected, this reduces exactly to the excited state mean-field theoryShea and Neuscamman (2018); Shea et al. (2020); Tran and Neuscamman (2020); Hardikar and Neuscamman (2020).
@@ II.1.2 Treatment of other excited states: double CIS
[…]
@@ II.2.1 Energy, gradients, and Hessian
It should first be noted that, for multistate calculations of standard MCSCF (without $\hat{P}$) such as SA-CASSCF, one frequently employs the following parameterization:Helgaker et al. (2000)
@@
The second derivatives with respect to $\bm{\lambda}$} are important not only for orbital optimization but also for geometry optimization (nuclear gradients), as will be discussed in a separate paper.Tsuchimochi (2026) Again, in real orbitals,
@@
As a useful interpretation of the orbital-optimization framework, it is worth mentioning that, for an RHF reference, SACIS can be interpreted as a special case of state-averaged RASSCF.Olsen et al. (1988); Malmqvist et al. (1990) Specifically, the RAS1 space corresponds to the full occupied orbital space, RAS2 is empty, and RAS3 contains the full virtual orbital space, with the restriction that at most one hole in RAS1 and one particle in RAS3 are allowed. In this sense, SACIS is formally equivalent to a state-averaged RASSCF description restricted to single excitations. Despite this formal connection, the SACIS implementation differs substantially from a conventional RASSCF algorithm. In particular, SACIS avoids molecular-orbital transformations of two-electron integrals by performing all contractions in the atomic-orbital basis, and it eliminates redundant orbital rotations within the occupied and virtual subspaces, which do not affect CIS energies. These features lead to a considerably more efficient implementation.
@@ III.1 Newton method and trust-region augmented Hessian method
However, this update is often not optimal when the states are (on average) far from the minimum, where the energy functional is not well approximated by a quadratic form.Helgaker et al. (2000) Instead, the trust-region augmented Hessian (TRAH) algorithm diagonalizes the augmented Hessian,
@@ III.3 Point-group symmetry
In spin-projection approaches, point-group symmetry is not explicitly enforced, and the converged wave function is therefore not guaranteed to retain the correct point-group symmetry, since spin-unrestricted methods allow the breaking of spatial symmetry in addition to spin symmetry. To explicitly account for point-group symmetry within the framework of symmetry breaking and restoration, a point-group symmetry projection may be desirable.Jiménez-Hoyos et al. (2012, 2013); Uejima and Ten-no (2017) However, from our experiences in many cases this is not needed; spin-projection can resurrect the lost spatial symmetry if not all. This means each of the contaminated spin states possesses a different point-group symmetry than the one the designated spin state has. Therefore, we will not perform the point-group symmetry projection, but simply assign each excited state to the dominant point-group irreducible representation (irrep).
@@ IV.1 Computational details
All CIS, ECIS, SACIS, SAECIS, DCIS, and EDCIS calculations were performed using the Gellan suite of programs.1 For complete-active-space SCF (CASSCF) and related methods, we used ORCA.Neese (2025) The 6-31G basis set was used for hydrogen fluoride and nitrogen, while aug-cc-pVDZ was employed for all other molecules.
@@
In state-averaged schemes, the number of states included in the averaging must be chosen appropriately. In practice, it is typically sufficient to include the lowest few states that may become nearly degenerate along the potential energy surface of interest. We find that the qualitative features of the results are not overly sensitive to moderate variations in the number of averaged states, provided that the relevant low-lying states are included. For the benchmark calculations in Section IV.3, we state-averaged over a fixed manifold of $n$ states that includes prescribed numbers of roots from each irrep in the benchmark set (including the ground state). In all methods, however, point-group symmetry constraints were not enforced during the orbital optimization. Therefore, for the SA optimization of multiple states with designated symmetry, we employed DIIS and, in each macro iteration, first converged more than $n$ CIS/ECIS roots in the micro step. From these converged roots, we then identified and selected the lowest-energy solutions with the desired irreps following the protocol described in Section III.3, and used only this selected set to construct the SA orbital gradient for the subsequent DIIS update.
@@ IV.3 Benchmark calculations for excitation energies of weakly correlated systems
Finally, we emphasize that the conclusions drawn here are specific to weakly correlated regimes. We expect that the proposed methods, particularly SAECIS and EDCIS, will perform more favorably for strongly correlated systems, where orbital relaxation and near-degeneracy effects play a more prominent role. A systematic assessment in such regimes, however, requires appropriate benchmark datasets, which are currently unavailable and will therefore be pursued in future work. In the absence of such a dataset, next we use the bond dissociation of hydrogen fluoride and nitrogen as stringent qualitative test cases for the proposed methods in strongly correlated regimes.
@@ IV.4 Potential energy curves of hydrogen fluoride
where $|\rm(Core)\rangle=|\phi_{{\rm F}_{1s}}^{2}\phi_{{\rm F}_{2s}}^{2}\phi_{{\rm F}_{p_{x}}}^{2}\phi_{{\rm F}_{p_{y}}}^{2}\rangle$ denotes the closed-shell, non-bonding orbitals. Evidence of bond breaking in orbital-optimized CIS was first discussed by Burton,Burton (2022) who studied the hydrogen molecule. Here we show that SACIS can also describe bond dissociation in hydrogen fluoride. Inspection of the ground-state wave function at large bond distances reveals that SACIS instead favors a localized-orbital description reminiscent of valence bond theory. In this limit, the reference determinant $|\Phi_{0}\rangle$ corresponds to the ionic configuration H+F-. Single excitations from $\phi_{{\rm F}_{p_{z}}}$ to $\phi_{{\rm H}_{1s}}$ within CIS then yield the correct dissociation limit in the form of a charge-transfer singlet,
@@
Consistent with this interpretation, the potential energy curve obtained with SACIS exhibits improved behavior near dissociation. As shown in Figure 4(b), the absolute ground-state energy errors of SACIS relative to FCI are comparable to those of SUHF. However, the SACIS curve is more nearly parallel to the FCI reference. This is quantified by the non-parallelity error (NPE), defined as the difference between the maximum and minimum deviations from FCI, which is reduced from 13.9 mHartree for SUHF to 9.6 mHartree for SACIS, as summarized in Table 3.
@@
Figure 5: Potential energy curves of N2 for the $S_{0}$$S_{3}$ states computed by different methods. All curves are shifted relative to the $S_{0}$ minimum at 1.1 Å.
IV.5 Potential energy curves of nitrogen
As a more stringent test case for strong correlation, we consider the ground and excited states of the nitrogen molecule at varying bond distances. In addition to the methods used in the preceding sections, we also include the MCTDA approach, i.e., the linear-response formulation of CASSCF within the TDA framework.Jørgensen et al. (1988) This allows us to examine how linear-response approaches, such as EDCIS and MCTDA, perform in describing excited states in the strongly correlated regime.
Figure 5 summarizes the potential energy curves of the ground state ($S_{0}$) and the three lowest singlet states ($S_{1}$$S_{3}$) regardless of irreducible representation. All curves are plotted relative to the $S_{0}$ energy at equilibrium ($R=1.1$ Å). Figure 5 clearly demonstrates that SACIS is not capable of describing the triple-bond dissociation without spin-symmetry breaking; its curves are qualitatively incorrect due to the lack of static correlation, showing a behavior similar to HF/CIS.
As expected, SSECIS shows qualitative agreement for $S_{0}$ with CASSCF and FCI. Its excitation energies computed with the linear-response EDCIS, however, fail to capture the strong correlation in the excited states near the dissociation limit. This behavior is similar to ECIS, where the ground state is described reasonably well but the excited states become inaccurate. This contrasts sharply with MCTDA, which shows remarkably similar results to FCI, even though these approaches are conceptually related in that a linear response is applied to the reference wave function (SSECIS or CASSCF(6e,6o)). The key difference is that MCTDA also includes the subspace spanned by all configuration variations within the active space. Consequently, the resulting excitation manifold generally contains higher excitation effects within the active space (e.g., configurations corresponding to double excitations relative to the ground-state CASSCF reference). In contrast, SSECIS does not employ an active space; therefore, the excitation manifold of EDCIS is spanned purely by single excitations from the SSECIS reference. This difference becomes particularly important in strongly correlated regimes, where single excitations alone cannot describe the multi-configurational character of excited states.
By contrast, the description provided by SAECIS is qualitatively acceptable not only for the $S_{0}$ state but also for the excited states across the dissociation region. Although genuine higher excitations are not explicitly included in SAECIS, their effects are partially recovered through orbital relaxation in the state-averaged framework (see the SACIS description of HF dissociation in the previous section) and through spin projection, which effectively introduces multi-configurational character into the reference wave function. This mechanism allows SAECIS to capture part of the static correlation that would otherwise require higher excitations in conventional CI-type methods.
The SAECIS curves agree well with those of SA-CASSCF(6e,6o) at shorter bond lengths (<2.0<2.0 Å), although the $S_{1}$$S_{3}$ energies begin to split at larger separations.At the dissociation limit ($R=3.0$ Å), the excitation energies of $S_{1}$, corresponding to the $\sigma\rightarrow\sigma^{*}$ transition, are 5.92, 5.64, 5.74, and 5.53 eV for SAECIS, SA-CASSCF(6e,6o), MCTDA, and FCI, respectively, showing good overall agreement. In contrast, the $S_{2}$ and $S_{3}$ states, corresponding to the $\sigma\rightarrow\pi^{*}$ transitions, lie 0.89 eV above $S_{1}$ with SAECIS, whereas the other methods, including ECIS, predict these states to be nearly degenerate with $S_{1}$. This behavior highlights a limitation of SAECIS in capturing the full degeneracy pattern in the strongly correlated dissociation limit; nevertheless, it still provides the most reasonable overall description among the approaches proposed in this work.